OpenStudy (anonymous):

How many solutions are there to the following system of equations? 4x – 14y = 6 –2x + 7y = –3 A. infinitely many B. 1 C. 2 D. 0

OpenStudy (anonymous):

@Hannah28

OpenStudy (anonymous):

Please help!!!!

OpenStudy (cwrw238):

multiply every term of second equation by -2 and see what you get

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

Can you tell me what infinetley many means?

OpenStudy (cwrw238):

–2x + 7y = –3 *-2 *-2 *-2 infinite means many many solutions with no end to the list

OpenStudy (anonymous):

so -4x 14y and -6?

OpenStudy (cwrw238):

no - not quite: - * - = + and - * + = - its 4x - 14y = 6 - the same as the first equation

OpenStudy (anonymous):

So then what do I do?

OpenStudy (cwrw238):

well there are many (infinite number) of values of x and y which fit this equatio

OpenStudy (anonymous):

oh ok thanks, could you help me on another?

OpenStudy (cwrw238):

there is only one equation not 2 and the graph of this equation is a straight line 

OpenStudy (cwrw238):

9x.y)as you can see there are an infinity number of points (x,y) on this line

OpenStudy (anonymous):

so wait the answer would be A. since you could put many diff numbers in

OpenStudy (cwrw238):

exactly

OpenStudy (cwrw238):

- well not just many - an infinite number

OpenStudy (anonymous):

Thanks can You help me in another one if I tag you in it?

OpenStudy (cwrw238):

sorry gotta go right away

OpenStudy (anonymous):

ok thanks for the help!!!

OpenStudy (anonymous):

@Lex_corney22 I'm sorry I was taking a test! At least you got help :)